OpenStudy (anonymous):

Help please solve the equation 3(c-2)=2(c-6) 2 -6 No Solution Identity

OpenStudy (john_es):

3c−6=2c−12⇒3c−2c=−12+6⇒c=-6

OpenStudy (anonymous):

THANK YOU